मराठी
तामिळनाडू बोर्ड ऑफ सेकेंडरी एज्युकेशनएचएससी विज्ञान इयत्ता १२

Choose the correct alternative: Consider a game where the player tosses a six-sided fair die. If the face that comes up is 6, the player wins ₹ 36, otherwise he lose

Advertisements
Advertisements

प्रश्न

Choose the correct alternative:

Consider a game where the player tosses a six-sided fair die. If the face that comes up is 6, the player wins ₹ 36, otherwise he loses ₹ k2, where k is the face that comes up k = {1, 2, 3, 4, 5}. The expected amount to win at this game in ₹ is

पर्याय

  • `19/6`

  • `- 19/6`

  • `3/2`

  • `- 3/2`

MCQ
Advertisements

उत्तर

`- 19/6`

shaalaa.com
Mathematical Expectation
  या प्रश्नात किंवा उत्तरात काही त्रुटी आहे का?
पाठ 11: Probability Distributions - Exercise 11.6 [पृष्ठ २१९]

APPEARS IN

सामाचीर कलवी Mathematics - Volume 1 and 2 [English] Class 12 TN Board
पाठ 11 Probability Distributions
Exercise 11.6 | Q 3 | पृष्ठ २१९

संबंधित प्रश्‍न

For the random variable X with the given probability mass function as below, find the mean and variance.

`f(x) = {{:(1/10, x = 2","  5),(1/5, x = 0","  1","  3","  4):}`


For the random variable X with the given probability mass function as below, find the mean and variance.

`f(x) = {{:(1/2 "e"^(x/2),  "for"  x > 0),(0,  "otherwise"):}`


If µ and σ2 are the mean and variance of the discrete random variable X and E(X + 3) = 10 and E(X + 3)2 = 116, find µ and σ2 


A lottery with 600 tickets gives one prize of ₹ 200, four prizes of ₹ 100, and six prizes of ₹ 50. If the ticket costs is ₹ 2, find the expected winning amount of a ticket


Choose the correct alternative:

A random variable X has binomial distribution with n = 25 and p = 0.8 then standard deviation of X is


Choose the correct alternative:

If P(X = 0) = 1 – P(X = 1). If E[X] = 3 Var(X), then P(X = 0) is


Find the expected value for the random variable of an unbiased die


Let X be a random variable defining number of students getting A grade. Find the expected value of X from the given table:

 X = x 0 1 2 3
P(X = x) 0.2 0.1 0.4 0.3

In a business venture a man can make a profit of ₹ 2,000 with a probability of 0.4 or have a loss of ₹ 1,000 with a probability of 0.6. What is his expected, variance and standard deviation of profit?


A person tosses a coin and is to receive ₹ 4 for a head and is to pay ₹ 2 for a tail. Find the expectation and variance of his gains


Let X be a random variable and Y = 2X + 1. What is the variance of Y if variance of X is 5?


Choose the correct alternative:

Which of the following is not possible in probability distribution?


Choose the correct alternative:

A probability density function may be represented by


Choose the correct alternative:

E[X – E(X)] is equal to


Choose the correct alternative: 

`int_(-oo)^oo` f(x) dx is always equal to


Choose the correct alternative: 

An expected value of a random variable is equal to it’s


What is the expected value of a game that works as follows: I flip a coin and if tails pay you ₹ 2; if heads pay you ₹ 1. In either case, I also pay you ₹ 0.50


Prove that V(X + b) = V(X)


The time to failure in thousands of hours of an important piece of electronic equipment used in a manufactured DVD player has the density function
f(x) = `{{:(2"e"^(-2x)",",  x > 0),(0",",  "otherwise"):}`
Find the expected life of this piece of equipment


Share
Notifications

Englishहिंदीमराठी


      Forgot password?
Use app×